OpenStudy (anonymous):

If you have 240 meters of fencing and want to enclose a rectangular area up against a long, straight wall, what is the largest area you can enclose?

OpenStudy (amistre64):

3x + y = 240 xy is a max use lagrange multipliers

OpenStudy (amistre64):

well, 2x + y = 240, thought it was a different problem

OpenStudy (anonymous):

It doesn't have to be that complicated I think... Think what kind of rectangle will give you the biggest area.

OpenStudy (amistre64):

granted there are other ways to approach it :)

OpenStudy (anonymous):

and by lagrange multipliers are you talking about the thing that you learn in calculus? O.o

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

What are those?

OpenStudy (anonymous):

Eh this is a kind of a question you would get in a grade 9 math xD

OpenStudy (anonymous):

Thats funny considering my calc 1 class is asking it

OpenStudy (amistre64):

if memory serves :) f(x,y) = xy g(x,y) = 2x + y - 240 fx = y, Lgx = 2L fy = x, Lgy = L x = L, y = 2L 2(L) + 3(2L) = 240 8L = 240, L=30 x = 30, y = 60

OpenStudy (amistre64):

typoed in the middle ...

OpenStudy (anonymous):

Thank you got it :)!

OpenStudy (amistre64):

2x + y = 240 2(L) + (2L) = 240 L = 60

OpenStudy (anonymous):

I dont know lol I got that question in my grade 9 math test lol I remember

OpenStudy (amistre64):

letting the area be defined as: xy and the constraint as: 2x+y = 240 we see from the constraint that y = 240-2x x(240-2x) is an upside down parabola with the highest point at x=60

OpenStudy (amistre64):

which would be the algebra way .... :)

OpenStudy (anonymous):

Let x and y be the sides of rectangle, then 2x+y=240 y=240-2x \[Area A=xy=x(240-2x)=240x-2x ^{2}\] \[\frac{ dA }{ dx }=240-4x\] \[\frac{ dA }{ dx }=0,240-4x=0,4x=240,x=\frac{ 240 }{ 4}=60\] \[\frac{ d ^{2}A }{ dx ^{2} }=-4\] \[at x=60, \frac{ d ^{2}A }{dx ^{2} }=-4<0\] Hence A is maximum at x=60 and y=240-2*60=240-120=120 Hence it is asquare with each side=120 m

OpenStudy (amistre64):

not a square